# Learning Decision-Focused Uncertainty Sets in Robust Optimization

### v4

@@ Appendix A Uncertainty sets and general reformulations
The Mean Robust Uncertainty (MRO) (Wang et al.,, 2022) set is defined as
${\mathcal{U}}_{\text{mro}}(\theta)$ =$=\left\{u=A(v_{11},\dots,v_{LK})+b\in S^{K\times L}\quad\middle|\quad\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}\|v_{lk}-\bar{d}_{k}\|^{p}\leq\rho^{p},\quad\alpha\in\Gamma\right\},$
where $\Gamma=\{\alpha~{}|~{}\sum_{l=1}^{L}\alpha_{lk}=w_{k},\alpha_{lk}\geq 0~{}\forall k,l\}$. Here, we assume the constraint function $g$ of the RO to be a maximum of $L$ pieces, indexed by $l$. We assume A is square, with dimensions ${\mbox{\bf R}}^{m\times m}$, and we have $\theta=(A,b)$. This set depends on $U^{N}$, a dataset of $N$ samples. We partition ${U}_{N}$ into $K$ disjoint subsets, with $\bar{d}_{k}$ the centroid of the $k$th subset. The weight we place upon each subset is $w_{k}$, and is equivalent to the proportion of points in the subset. We choose $p$ to be an integer exponent, and $\rho$ to be a positive radius. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&\lambda\rho^{p}+\sum_{k=1}^{K}w_{k}s_{k}\leq 0\\ &[-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}\\ &\qquad\qquad+\phi(q)\lambda\left\|A^{T}(h_{lk}+\omega_{lk})/\lambda\right\|^{q}_{*}\leq s_{k},\quad l=1,\dots,L,\quad k=1,\dots,K\\ &\lambda\geq 0.\end{array}$
with auxiliary variables $s,h_{lk},\omega_{lk}\in{\mbox{\bf R}}^{m}$ for $j=1,\dots,J$, $k=1,\dots,K$, and $\lambda\in{\mbox{\bf R}}$. For the special case of the maximum of affine constraint, we have
minimize $\begin{array}[]{ll}\text{minimize }&f(x)\\ \mbox{subject to}&\lambda\rho^{p}+\sum_{k=1}^{K}w_{k}s_{k}\leq 0\\ &a_{l}(x)+\sigma_{S}(\omega_{lk})-b^{T}(\omega_{lk}-P_{l}(x))-(A^{T}(\omega_{lk}-P_{l}(x)))^{T}\bar{d}_{k}\\ &\qquad\qquad+\phi(r)\lambda\left\|A^{T}(\omega_{lk}-P_{l}(x))/\lambda\right\|^{r}_{q}\leq s_{k},\quad l=1,\dots,L,\quad k=1,\dots,K\\ &\lambda\geq 0.\end{array}$ (20)
@@ A.1 Linear transformation to cone programs in standard form
The convex optimization problems above can be transformed to cone programs (Agrawal et al., 2019a, ) in the form
$\begin{array}[]{ll}\mbox{minimize}&c^{T}x\\ \mbox{subject to}&Gx+s=h\\ &s\in\mathcal{K},\end{array}$ (21)
where $x\in{\mbox{\bf R}}^{\bar{n}}$ is the primal variable, $s\in{\mbox{\bf R}}^{\bar{m}}$ is a slack variable, and the set $\mathcal{K}\subseteq{\mbox{\bf R}}^{\bar{m}}$ is a nonempty, closed, convex cone with dual cone $\mathcal{K}^{\star}\subseteq{\mbox{\bf R}}^{\bar{m}}$. So long as the constituent functions $g_{l}$ of the constraint $g$ from the robust problem (1) are concave in the uncertain parameters, and that these functions satisfy the (disciplined parametrized program) DPP rules proposed by Agrawal et al., 2019a , the cone program’s data $(G,h,c)$ can be expressed as a linear function (i.e., tensor product) of the parameters $\theta$ and $y$ (Agrawal et al., 2019a, , Lemma 1). The DPP rules depend on well-known composition theorems for convex functions, and ensure that every function in the program is convex in the decision variables, or affine if representing equalities.
@@ Appendix B Conservative Jacobians for the augmented Lagrangian
We give the exact forms of the conservative Jacobians from Proposition 3.1. In this section, for simplicity, we denote the solution of the optimization problem (1), $x^{j}=x(\theta,y^{j})$, and its conservative Jacobian, $J_{x^{j}}(\theta)$, for each input pair $(\theta,y^{j})$. For the Lagrangian function $\hat{L}(\alpha,s,\theta,\lambda,\mu)$ over the batch $B=(U\times Y)\subseteq(U^{N}\times Y^{J})$, where $|B|$ denotes the cardinality of the set, we have the following conservative Jacobians.
$J_{\hat{L}}(\theta)$ $=\frac{1}{|Y|}\sum_{y^{j}\in Y}J_{f}(x^{j})J_{x^{j}}(\theta)$
$+\frac{\lambda}{\eta|B|}\sum_{(u^{i},y^{j})\in B}\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha\right\}J_{g}(x^{j})J_{x^{j}}(\theta)$
$+\frac{\mu}{\eta|B|}\sum_{(u^{i},y^{j})\in B}((g(x^{j},u^{i},y^{j})-\alpha)/\eta+\alpha+s-\kappa)\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha\right\}J_{g}(x^{j})J_{x^{j}}(\theta),$
$J_{\hat{L}}(s)$ $=\lambda+\frac{\mu}{|B|}\sum_{(u^{i},y^{j})\in B}(g(x^{j},u^{i},y^{j})-\alpha)_{+}/\eta+\alpha+s-\kappa,$
$J_{\hat{L}}(\alpha)$ $=\frac{\lambda}{|B|}\sum_{(u^{i},y^{j})\in B}(1-(1/\eta)\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha\right\})$
$+\frac{\mu}{|B|}\sum_{(u^{i},y^{j})\in B}(1-(1/\eta)\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha\right\})((g(x^{j},u^{i},y^{j})-\alpha)_{+}/\eta+\alpha+s-\kappa).$
For a condition $A$, we denote its indicator function $\mathbf{1}\{A\}$, taking the value 1 when true and 0 when false. When the Lagrangian function is defined over the entire dataset, we formulate the conservative Jabocian $J_{L}(\alpha,s,\theta)$ the same way as above, with $B=(U^{N}\times Y^{J})$.
@@ Appendix C Proof of conservative Jacobians for the robust optimization problem
To prove Theorem 3.1, we characterize the path-differentiability of the function $x(\theta,y)$ by solving the inner optimization problem and obtaining conservative Jacobians using the results of Agrawal et al., 2019b and Bolte et al., (2021). The optimization problem (1) is first reformulated into a convex and tractable problem in conic form (21), as shown in Appendix A.1. The conic program’s data, $(G,h,c)$, is parameterized by $(\theta,y)$, through the mapping
$(\theta,y)\rightarrow(G,h,c),$
which is linear in $\theta$ and $y$ for all uncertainty sets in Appendix A, given the disciplined parametrized program (DPP) requirement detailed in Appendix A.1 (Agrawal et al., 2019a, , Lemma 1).
This formulation has a corresponding dual, with dual variable $r\in{\mbox{\bf R}}^{m}$, and solving the conic program amounts to solving the homogeneous self-dual embedding (Ye et al.,, 1994) of the primal-dual problem. In the following, we define $N=\bar{n}+\bar{m}+1$. The overall procedure is split into three steps: mapping the problem data ($G,h,c$) onto a skew symmetric matrix $Q$:
$Q=\left[\begin{array}[]{lll}0&G^{T}&c\\ -G&0&h\\ -c^{T}&-h^{T}&0\end{array}\right]\in{\mbox{\bf R}}^{N\times N},$
finding the solution $z=(u,v,w)\in{\mbox{\bf R}}^{N}$ to the homogeneous self-dual embedding, and mapping $z$ onto the solution of the primal-dual pair (Agrawal et al., 2019b, ). We can express the chain of mappings as
$(G,h,c)\rightarrow Q\rightarrow z\rightarrow(x,r,s).$
To solve for $z=\nu(Q)$, shown as the mapping $Q\rightarrow z$, we define the normalized residual map (Busseti et al.,, 2019), given as
$\mathcal{N}(z,Q)=((Q-I)\Pi+I)(z/|w|),$
where $\Pi$ is the projection onto the convex cone $\mathcal{K}^{\prime}={\mbox{\bf R}}^{\bar{n}}\times\mathcal{K}^{\star}\times{\mbox{\bf R}}_{+}$. For a given $Q$, $z=\nu(Q)$ is a solu tion to the homogeneous self-dual embedding if and only if $\mathcal{N}(z,Q)=0$ and $w>0$. Once this solution is obtained, the mapping $z\rightarrow(x,r,s)$ is given by the function
$\phi(z)=(u,\Pi(v),\Pi(v)-v)/w.$
We can now invoke proposition 4 of the work by Bolte et al., (2021), together with the results from Agrawal et al., 2019b to show that $\nu$ and $\phi$ are path-differentiable.
@@
Assume the functions $\Pi$ and $\mathcal{N}$ are path differentiable, with conservative Jacobians $J_{\Pi}$ and $J_{\mathcal{N}}$. Assume, for all $Q$ constructed from the problem data $(G,h,c)$, and $z=\nu(Q)$ is the unique solution to $\mathcal{N}(z,Q)=0$. Furthermore, assume all matrices formed by the first $N$ columns of $J_{\mathcal{N}}(z,Q)$ are invertible. Then, the functions $\nu$ and $\phi$ are path-differentiable, with conservative Jacobians
$J_{\nu}(Q)=\{-U^{-1}V\mid[U~{}V]\in J_{\mathcal{N}}(z,Q)\},\qquad J_{\phi}(z)=\left[\begin{array}[]{ccc}I&0&-x\\ 0&J_{\Pi}(v)&-r\\ 0&J_{\Pi}(v)-I&-s\end{array}\right].$
The assumptions on $\Pi$ and $\mathcal{N}$ hold, as the cones we consider (zero cones, free cones, nonnegative cones, and second-order cones) are sub-differentiable (Busseti et al.,, 2019; Kanzow et al.,, 2009). By Theorem 1 of the work by Bolte et al., (2021), which states that the Clarke Jacobian is a minimal conservative Jacobian, we conclude that the conic projections admit conservative Jacobians, and are thus path differentiable. It follows that $\mathcal{N}$, as a linear function of $\Pi$ and $Q$, is also path-differentiable. The invertibility of $U$ holds as $U=((Q-I)J_{\Pi}(z)+I)/w$ (Agrawal et al., 2019b, ), which is nonsingular due to the PSD nature of the members of the set $J_{\Pi}(z)$. Lastly, the assumption of uniqueness holds from Assumption 3.3.
@@
As $(x,r,s)=\phi(z)$, we can partition $J_{\phi}(\theta)$ as $(J_{x}(\theta),J_{r}(\theta),J_{s}(\theta))$ where the first component is the conservative Jacobian of interest. Therefore, we conclude that $x(\theta,y)$ is path differentiable, with nonempty conservative Jacobian $J_{x}(\theta)$.
@@ Appendix D Proof of convergence of the augmented Lagrangian algorithm
We prove Theorem 3.2, the convergence of the augmented Lagrangian algorithm to a stationary point. We separate the procedure into three steps. In the first step, we show for Algorithm 2 with fixed $\lambda,\mu$, the convergence of $\{\alpha^{t},s^{t},\theta^{t}\}$ to a stationary point of the augmented Lagrangian function (11) as $t\to\infty$. In particular, we note that by Theorem 1 of Bolte and Pauwels, (2019), conservative Jacobians are gradients almost everywhere, so by Bolte et al., (2021), the iterates of the algorithm is a Clarke stochastic subgradient sequence almost surely. We thus use the Clarke differential (Clarke,, 1975), $\partial^{c}f(x)\coloneqq{\rm conv}\{\lim_{x_{i}\rightarrow x}\nabla f(x_{i})\mid x_{i}\in{\rm dom}(\nabla f)\}$ to denote the subdifferentials in our analysis, instead of using $J_{f}(x)$. In the next two steps, we show for Algorithm 1, either the convergence of $\{\theta^{k}\}$ to a feasible point of conditon (9), or the convergence of $\{\alpha^{k},s^{k},\theta^{k}\}$ to a stationary point of the constraint violation function.
Step 1: Convergence of ($\alpha$, $s$, $\theta$).
@@
In this step, we consider the augmented Lagrangian function with fixed $\lambda$ and $\mu$. By Assumptions 3.4 and 3.5 on the boundedness and step sizes of the iterates, by the definability of the augmented Lagrangian function, and by the existence of its conservative Jacobians, we apply Theorem 3 of Bolte et al., (2021), on the convergence of SGD for path-differentiable functions. The projection onto the nonnegative orthant for $s$ does not affect the convergence results, as proximal SGD with conservative Jacobians also converges, subject to the same assumptions (Davis et al.,, 2018, Theorem 6.2). Let $L(\alpha,s,\theta)=L(\alpha,s,\theta,\lambda,\mu)$, as the suppressed variables are fixed, and let $(\alpha^{\star},s^{\star},\theta^{\star})$ be a local minimum point of this function. By the aforementioned theorems, we can show that $V_{\lambda,\mu}(\alpha,s,\theta)\coloneqq L(\alpha,s,\theta)-L(\alpha^{\star},s^{\star},\theta^{\star})$ is a Lyapunov function for the differential inclusion $(\dot{\alpha},\dot{s},\dot{\theta})\in-\partial^{c}L(\alpha,s,\theta)$. Specifically, there exist a ball of radius $r$ around $(\alpha^{\star},s^{\star},\theta^{\star})$ such that for $(\alpha,s,\theta)\in\mathbf{B}_{\alpha^{\star},s^{\star},\theta^{\star}}(r)$, the Lyapunov function is locally positive definite. This function is decreasing along solutions strictly outside of the set of Clarke-critical points $\{(\alpha,s,\theta)\in{\mbox{\bf R}}^{p+2},0\in\partial^{c}L(\alpha,s,\theta)\}$. Then, for almost all initializations $(\alpha^{0},s^{0},\theta^{0})\in\mathcal{D}_{\alpha}\times\mathcal{D}_{s}\times\mathcal{D}_{\theta}$, the objective $L(\alpha^{t},s^{t},\theta^{t})$ converges almost surely and all stationary points $(\bar{\alpha},\bar{s},\bar{\theta})$ of $\{\alpha^{t},s^{t},\theta^{t}\}$ are Clarke-critical in the sense that $0\in\partial^{c}L(\bar{\alpha},\bar{s},\bar{\theta})$ (Bolte et al.,, 2021). We thus conclude that, with almost any initial point $(\alpha^{0},s^{0},\theta^{0})\in\mathbf{B}_{\alpha^{\star},s^{\star},\theta^{\star}}(r)\subseteq\mathcal{D}_{\alpha}\times\mathcal{D}_{s}\times\mathcal{D}_{\theta}$, the sequence $\{\alpha^{t},s^{t},\theta^{t}\}$ converges to a stationary point $(\alpha^{\star},s^{\star},\theta^{\star})$, satisfying $0\in J_{L}(\alpha^{\star},s^{\star},\theta^{\star})$.

### v5

@@ Appendix A Uncertainty sets and general reformulations
The Mean Robust Uncertainty (MRO) (Wang et al.,, 2024) set is defined as
${\mathcal{U}}_{\text{mro}}(\theta)$ =$=\left\{u=A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}(v_{11},\dots,v_{LK})+b\in S^{K\times L}\quad\middle|\quad\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}\|v_{lk}-\bar{d}_{k}\|^{p}\leq\rho^{p},\quad\alpha\in\Gamma\right\},$
where $\Gamma=\{\alpha~{}|~{}\sum_{l=1}^{L}\alpha_{lk}=w_{k},\alpha_{lk}\geq 0~{}\forall k,l\}$. Here, we assume the constraint function $g$ of the RO to be a maximum of $L$ pieces, indexed by $l$. We assume A is square, with dimensions ${\mathbf{R}}^{m\times m}$, and we have $\theta=(A(\cdot),b(\cdot))$. This set depends on $U^{N}$, a dataset of $N$ samples. We partition ${U}_{N}$ into $K$ disjoint subsets, with $\bar{d}_{k}$ the centroid of the $k$th subset. The weight we place upon each subset is $w_{k}$, and is equivalent to the proportion of points in the subset. We choose $p$ to be an integer exponent, and $\rho$ to be a positive radius. For the maximum of concave constraint, we have the following reformulation,
$\begin{array}[]{ll}{\text{minimize}}&f(z)\\ \mbox{subject to}&\lambda\rho^{p}+\sum_{k=1}^{K}w_{k}s_{k}\leq 0\\ &[-g_{l}]^{\star}(x,h_{lk})-b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(h_{lk}+\omega_{lk})+\sigma_{S}(\omega_{lk})-(A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}\\ &\qquad\qquad+\phi(q)\lambda\left\|A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(h_{lk}+\omega_{lk})/\lambda\right\|^{q}_{*}\leq s_{k},\quad l=1,\dots,L,\quad k=1,\dots,K\\ &\lambda\geq 0.\end{array}$
with auxiliary variables $s,h_{lk},\omega_{lk}\in{\mathbf{R}}^{m}$ for $j=1,\dots,J$, $k=1,\dots,K$, and $\lambda\in{\mathbf{R}}$. For the special case of the maximum of affine constraint, we have
minimize $\begin{array}[]{ll}\text{minimize }&f(z)\\ \mbox{subject to}&\lambda\rho^{p}+\sum_{k=1}^{K}w_{k}s_{k}\leq 0\\ &a_{l}(z)+\sigma_{S}(\omega_{lk})-b{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(\omega_{lk}-P_{l}(x))-(A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(\omega_{lk}-P_{l}(z)))^{T}\bar{d}_{k}\\ &\qquad\qquad+\phi(r)\lambda\left\|A{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(x)}^{T}(\omega_{lk}-P_{l}(z))/\lambda\right\|^{r}_{q}\leq s_{k},\quad l=1,\dots,L,\quad k=1,\dots,K\\ &\lambda\geq 0.\end{array}$ (22)
@@ A.1 Linear transformation to cone programs in standard form
The convex optimization problems above can be transformed to the cone program (Agrawal et al., 2019a, )
$\begin{array}[]{ll}\mbox{minimize}&c^{T}z\\ \mbox{subject to}&Gz+s=h\\ &s\in\mathcal{K},\end{array}$ (23)
where $z\in{\mathbf{R}}^{\bar{n}}$ is the primal variable, $s\in{\mathbf{R}}^{\bar{m}}$ is a slack variable, and the set $\mathcal{K}\subseteq{\mathbf{R}}^{\bar{m}}$ is a nonempty, closed, convex cone with dual cone $\mathcal{K}^{\star}\subseteq{\mathbf{R}}^{\bar{m}}$. So long as the constituent functions $g_{l}$ of the constraint $g$ from the robust problem (1) are concave in the uncertain parameters, and that these functions satisfy the (disciplined parametrized program) DPP rules proposed by Agrawal et al., 2019a , the cone program’s data $(G,h,c)$ can be expressed as a linear function (i.e., tensor product) of the parameters $\theta$ and $x$ (Agrawal et al., 2019a, , Lemma 1). The DPP rules depend on well-known composition theorems for convex functions, and ensure that every function in the program is convex in the decision variables, or affine if representing equalities.
@@ Appendix B Conservative Jacobians for the augmented Lagrangian
We give the conservative Jacobians from Proposition 3.1. In this section, for simplicity, we denote the solution of the optimization problem (1), $z^{i}=z(\theta,x^{i})$, and its conservative Jacobian, $J_{z^{i}}(\theta)$, for each input pair $(\theta,x^{i})$. For the Lagrangian function $\hat{L}(\alpha,\theta,\lambda,\mu)$ over the batch $B=(U,X)\subseteq W^{N}$, where $|B|$ denotes the cardinality of the set, we have the following conservative Jacobians,
$J_{\hat{L}}(\theta)$ $=\frac{1}{|B|}\sum_{x^{i}\in Y}J_{\Phi}(z^{i})J_{z^{i}}(\theta)+\frac{\lambda}{\eta|B|}\sum_{(u^{i},x^{i})\in B}M^{i}\mathbf{1}\left\{g(z^{i},u^{i},x^{i})\geq\alpha\right\}J_{g}(z^{i})J_{z^{i}}(\theta)$
$+\frac{\mu}{\eta|B|}\sum_{(u^{i},x^{i})\in B}M^{i}((g(z^{i},u^{i},x^{i})-\alpha)/\eta+\alpha+s-\kappa)\mathbf{1}\left\{g(z^{i},u^{i},x^{i})\geq\alpha\right\}J_{g}(z^{i})J_{z^{i}}(\theta),$
$J_{\hat{L}}(\alpha)$ $=\frac{\lambda}{|B|}\sum_{(u^{i},x^{i})\in B}M^{i}(1-(1/\eta)\mathbf{1}\left\{g(z^{i},u^{i},x^{i})\geq\alpha\right\})$
$+\frac{\mu}{|B|}\sum_{(u^{i},x^{i})\in B}M^{i}(1-(1/\eta)\mathbf{1}\left\{g(z^{i},u^{i},x^{i})\geq\alpha\right\})((g(z^{i},u^{i},x^{i})-\alpha)_{+}/\eta+\alpha+s-\kappa),$
where $M^{i}=\mathbf{1}\left\{(g(z^{i},u^{i},x^{i})-\alpha^{k})_{+}/\eta+\alpha^{k}-\kappa\geq 0\right\}.$ For a condition $A$, we denote its indicator function $\mathbf{1}\{A\}$, taking the value 1 when true and 0 when false. When the Lagrangian function is defined over the entire dataset, we formulate the conservative Jabocian $J_{L}(\alpha,s,\theta)$ the same way as above, with $B=W^{N}$.
@@ Appendix C Proof of conservative Jacobians for the robust optimization problem
To prove Theorem 3.1, we characterize the path-differentiability of the function $z(\theta,x)$ by solving the inner optimization problem and obtaining conservative Jacobians using the results of Agrawal et al., 2019b and Bolte et al., (2021). The optimization problem (1) is first reformulated into a convex and tractable problem in conic form (23), as shown in Appendix A.1. The conic program’s data, $(G,h,c)$, is parameterized by $(\theta,x)$, through the mapping
$(\theta,x)\rightarrow(G,h,c),$
which is linear in $\theta$ and $x$ for all uncertainty sets in Appendix A, given the disciplined parametrized program (DPP) requirement detailed in Appendix A.1 (Agrawal et al., 2019a, , Lemma 1), as well as the linear contextual mapping we defined in Section 5.
This formulation has a corresponding dual, with dual variable $r\in{\mathbf{R}}^{m}$, and solving the conic program amounts to solving the homogeneous self-dual embedding (Ye et al.,, 1994) of the primal-dual problem. In the following, we define $N=\bar{n}+\bar{m}+1$. The overall procedure is split into three steps: mapping the problem data ($G,h,c$) onto a skew symmetric matrix $Q$:
$Q=\left[\begin{array}[]{lll}0&G^{T}&c\\ -G&0&h\\ -c^{T}&-h^{T}&0\end{array}\right]\in{\mathbf{R}}^{N\times N},$
finding the solution $y=(u,v,w)\in{\mathbf{R}}^{N}$ to the homogeneous self-dual embedding, and mapping $y$ onto the solution of the primal-dual pair (Agrawal et al., 2019b, ). We can express the chain of mappings as
$(G,h,c)\rightarrow Q\rightarrow y\rightarrow(z,r,s).$
To solve for $y=\nu(Q)$, shown as the mapping $Q\rightarrow y$, we define the normalized residual map (Busseti et al.,, 2019), given as
$\mathcal{N}(y,Q)=((Q-I)\Pi+I)(y/|w|),$
where $\Pi$ is the projection onto the convex cone $\mathcal{K}^{\prime}={\mathbf{R}}^{\bar{n}}\times\mathcal{K}^{\star}\times{\mathbf{R}}_{+}$. For a given $Q$, $y=\nu(Q)$ is a solution to the homogeneous self-dual embedding if and only if $\mathcal{N}(y,Q)=0$ and $w>0$. Once this solution is obtained, the mapping $y\rightarrow(z,r,s)$ is given by the function $\phi(y)=(u,\Pi(v),\Pi(v)-v)/w.$ We can now invoke proposition 4 of the work by Bolte et al., (2021), together with the results from Agrawal et al., 2019b to show that $\nu$ and $\phi$ are path-differentiable.
@@
Assume the functions $\Pi$ and $\mathcal{N}$ are path differentiable, with conservative Jacobians $J_{\Pi}$ and $J_{\mathcal{N}}$. Assume, for all $Q$ constructed from the problem data $(G,h,c)$, and $y=\nu(Q)$ is the unique solution to $\mathcal{N}(y,Q)=0$. Furthermore, assume all matrices formed by the first $N$ columns of $J_{\mathcal{N}}(y,Q)$ are invertible. Then, the functions $\nu$ and $\phi$ are path-differentiable, with conservative Jacobians
$J_{\nu}(Q)=\{-U^{-1}V\mid[U~{}V]\in J_{\mathcal{N}}(y,Q)\},\qquad J_{\phi}(y)=\left[\begin{array}[]{ccc}I&0&-z\\ 0&J_{\Pi}(v)&-r\\ 0&J_{\Pi}(v)-I&-s\end{array}\right].$
The assumptions on $\Pi$ and $\mathcal{N}$ hold, as the cones we consider (zero cones, free cones, nonnegative cones, and second-order cones) are sub-differentiable (Busseti et al.,, 2019; Kanzow et al.,, 2009). By Theorem 1 of the work by Bolte et al., (2021), which states that the Clarke Jacobian is a minimal conservative Jacobian, we conclude that the conic projections admit conservative Jacobians, and are thus path differentiable. It follows that $\mathcal{N}$, as a linear function of $\Pi$ and $Q$, is also path-differentiable. The invertibility of $U$ holds as $U=((Q-I)J_{\Pi}(y)+I)/w$ (Agrawal et al., 2019b, ), which is nonsingular due to the PSD nature of the members of the set $J_{\Pi}(y)$ and the skew-symmetric nature of $Q$, so long as $(Q-I)J_{\Pi}(y)$ does not have eigenvalues of $-1$. Lastly, the assumption of uniqueness holds from Assumption 3.3.
@@
As $(z,r,s)=\phi(y)$, we can partition $J_{\phi}(\theta)$ as $(J_{z}(\theta),J_{r}(\theta),J_{s}(\theta))$ where the first component is the conservative Jacobian of interest. Therefore, we conclude that $z(\theta,x)$ is path differentiable, with nonempty conservative Jacobian $J_{z}(\theta)$.
Lastly, we note that when the matrix $U$ is singular, we can still solve for the Jacobian using a least-squares approximation, as remarked in (Agrawal et al., 2019a, , Appendix B). In fact, for large matrices, the least-squares approximation is the most efficient solution method for all problems.
@@ Appendix D Proof of convergence of the augmented Lagrangian algorithm
We prove Theorem 3.2, the convergence of the augmented Lagrangian algorithm to a stationary point. We separate the procedure into three steps. In the first step, we show for Algorithm 2 with fixed $\lambda,\mu$, the convergence of $\{\alpha^{t},\theta^{t}\}$ to a stationary point of the augmented Lagrangian function (14) as $t\to\infty$. In particular, we note that by Theorem 1 of Bolte and Pauwels, (2019), conservative Jacobians are gradients almost everywhere, so by Bolte et al., (2021), the iterates of the algorithm is a Clarke stochastic subgradient sequence almost surely. We thus use the Clarke differential (Clarke,, 1975), $\partial^{c}f(z)\coloneqq{\rm conv}\{\lim_{z_{i}\rightarrow z}\nabla f(z_{i})\mid z_{i}\in{\rm dom}(\nabla f)\}$ to denote the subdifferentials in our analysis, instead of using $J_{f}(z)$. In the next two steps, we show for Algorithm 1, either the convergence of $\{\theta^{k}\}$ to a feasible point of condition (12), or the convergence of $\{\alpha^{k},\theta^{k}\}$ to a stationary point of the constraint violation function.
Step 1: Convergence of ($\alpha$, $\theta$).
@@
In this step, we consider the augmented Lagrangian function with fixed $\lambda$ and $\mu$. By Assumptions 3.4 and 3.5 on the boundedness and step sizes of the iterates, by the definability of the augmented Lagrangian function, and by the existence of its conservative Jacobians, we apply Theorem 3 of Bolte et al., (2021), on the convergence of SGD for path-differentiable functions. The projection onto the nonnegative orthant for $s$ does not affect the convergence results, as proximal SGD with conservative Jacobians also converges, subject to the same assumptions (Davis et al.,, 2018, Theorem 6.2). Let $L(\alpha,\theta)=L(\alpha,\theta,\lambda,\mu)$, as the suppressed variables are fixed, and let $(\alpha^{\star},\theta^{\star})$ be a local minimum point of this function. By the aforementioned theorems, we can show that $V_{\lambda,\mu}(\alpha,\theta)\coloneqq L(\alpha,\theta)-L(\alpha^{\star},\theta^{\star})$ is a Lyapunov function for the differential inclusion $(\dot{\alpha},\dot{\theta})\in-\partial^{c}L(\alpha,\theta)$. Specifically, there exist a ball of radius $r$ around $(\alpha^{\star},\theta^{\star})$ such that for $(\alpha,\theta)\in\mathbf{B}_{\alpha^{\star},\theta^{\star}}(r)$, the Lyapunov function is locally positive definite. This function is decreasing along solutions strictly outside of the set of Clarke-critical points $\{(\alpha,\theta)\in{\mathbf{R}}^{p+1},0\in\partial^{c}L(\alpha,\theta)\}$. Then, for almost all initializations $(\alpha^{0},\theta^{0})\in\mathcal{D}_{\alpha}\times\mathcal{D}_{\theta}$, the objective $L(\alpha^{t},\theta^{t})$ converges almost surely and all stationary points $(\bar{\alpha},\bar{\theta})$ of $\{\alpha^{t},\theta^{t}\}$ are Clarke-critical in the sense that $0\in\partial^{c}L(\bar{\alpha},\bar{\theta})$ (Bolte et al.,, 2021). We thus conclude that, with almost any initial point $(\alpha^{0},\theta^{0})\in\mathbf{B}_{\alpha^{\star},\theta^{\star}}(r)\subseteq\mathcal{D}_{\alpha}\times\mathcal{D}_{\theta}$, the sequence $\{\alpha^{t},\theta^{t}\}$ converges to a stationary point $(\alpha^{\star},\theta^{\star})$, satisfying $0\in J_{L}(\alpha^{\star},\theta^{\star})$.